# Accelerating Containerized Service Delivery at the Network Edge

### v1

@@ Abstract
Efficient container image distribution is crucial for enabling machine learning inference at the network edge, where resource limitations and dynamic network conditions create significant challenges. In this paper, we present PeerSync, a decentralized P2P-based system designed to optimize image distribution in edge environments. PeerSync employs a popularity- and network-aware download engine that dynamically adapts to content popularity and real-time network conditions using a sliding window mechanism. PeerSync further integrates automated tracker election for rapid peer discovery and dynamic cache management for efficient storage utilization. We implement PeerSync with 8000+ lines of Rust code and test its performance extensively on both physical edge devices and Docker-based emulations. Experimental results show that PeerSync delivers a remarkable speed increase of 2.72$\times$, 1.79$\times$, and 1.28$\times$ compared to the Baseline, Dragonfly, and Kraken, respectively, while significantly reducing peak cross-network traffic by 90.72% under congested and varying network conditions.
@@
edge computing, container image distribution, P2P architecture, model inference.
@@ I Introduction
Deploying a containerized application involves retrieving the necessary image layers from a container registry (e.g., Docker Hub), unpacking them, mounting them using a layered filesystem, and initiating the container entrypoint. These images, structured as a stack of layers representing filesystem segments, are stored and managed in centralized container registries. While these registries efficiently facilitate container image distribution in cloud environments, they encounter unique challenges in edge computing scenarios.
  • High latency. Distributed edge devices face substantial delays when fetching image layers over wide-area networks, leading to degraded responsiveness and increased cold-startup times for services [16, 8, 11].
  • Inefficiency under high concurrency. The non-parallel architecture of container registry uplinks often causes severe bottlenecks when handling concurrent image-pulling requests from multiple edge devices [11]. The growing size of AI models exacerbates this issue.
To address these limitations, peer-to-peer (P2P) architectures have been explored to accelerate container image distribution while alleviating the load on centralized registries [16, 44, 48, 19, 22, 26, 49, 23, 27, 2, 43]. P2P systems inherently offer distributed replication, parallel data transfers, and resilience against single points of failure. Although P2P-based solutions have shown promise in cloud environments, their effectiveness in edge settings is constrained by the unique characteristics of edge computing:
  • Limited network bandwidth. Unlike high-capacity data center networks, edge environments often operate with significantly constrained bandwidth, hindering efficient data exchange [13].
@@
  • Weak stability under dynamic topology. The dynamic nature of edge networks introduces frequent connectivity changes, requiring robust mechanisms to adaptively manage edge devices and maintain consistency without relying on centralized components such as static trackers [20]. [^1] A tracker is used to identify which peers possess or request a given object in a P2P network. See [20] for more details.
While P2P-based solutions such as EdgePier [8], Kraken [43], and Starlight [11] address some of these issues and achieve great performance in most cases, they are still based on the traditional P2P architecture and consequently fall short in effectively balancing network efficiency, storage optimization, and adaptability to edge-specific conditions. More specifically, they lack fine-grained control mechanisms and rely on static trackers, making them unsuitable for fluctuating network conditions. Starlight [11] minimizes container startup latency by redesigning deployment protocols and storage formats but requires intrusive modifications to cloud registries, limiting its applicability. Moreover, existing solutions often neglect factors such as dynamic network quality, content popularity, and effective storage utilization, leaving room for improvement. For instance, our experiments (detailed in Sec. II) demonstrate that maintaining a single copy of any image layer within a local area network (LAN) can significantly enhance image-fetching speed without overloading local storage.
TABLE I: A comparison between our work and state-of-the-art implementations. ‘F.D.’ stands for ‘Fully Decentralized’, while ‘S.D.’ stands for ‘Semi Decentralized’.
Reference PeerSync Kraken [43] Dragonfly [2] Starlight [11] EdgePier [8] FID [23]
@@
P2P flavor Impl. from scratch Impl. from scratch Impl. from scratch N/A IPFS BitTorrent
Scenario Edge Cloud Cloud WAN Edge General
Tracker placement dynamic static static N/A static static
Open-sourced
To overcome these challenges, we propose PeerSync, a non-intrusive, P2P-based system tailored to the unique demands of edge computing. PeerSync constructs a fully decentralized P2P network spanning multiple edge devices across different LANs. It features a P2P image download engine that optimizes distribution by periodically calculating the required content pieces for each missing layer of a requested image, leveraging a scoring system that evaluates both content popularity and the quality of the network connection between peers and the requesting edge device. Unlike traditional P2P architectures that rely on manually configured trackers, PeerSync autonomously elects trackers based on real-time metrics, including bandwidth availability and resource utilization, ensuring resilience and performance. PeerSync also employs a selective deletion mechanism to optimize storage utilization, maintaining only essential image layers within each LAN. We compare PeerSync with state-of-the-art solutions in Table I. In summary, our main contributions are as follows:
@@
    We accelerate edge container image distribution with a tailored P2P architecture, featuring a high-performance download engine, fault-tolerant tracker with election ability for faster peer discovery, and collaborative dynamic space reclamation for efficient storage management.
@@
    We dynamically assign scores to peers based on runtime metrics like content popularity and network conditions using a sliding window mechanism, significantly enhancing the flexibility of image distribution in dynamic edge networks.
@@
    We implement PeerSync with 8000+ lines of Rust code, resulting in a statically linked binary of just 8.8 MB, making it easily deployable on edge devices.
@@
    We conducted extensive experiments to verify PeerSync’s performance under different network conditions. On average, PeerSync achieves 2.72x faster distribution than the Baseline, 1.79x faster than Dragonfly, and 1.28x faster than Kraken. Additionally, PeerSync reduces peak cross-network traffic by 90.72% under congested and variable network conditions.
@@ II Motivation
In this section, we demonstrate the limitations of existing approaches using Kraken [43] and Starlight [11], motivating the need for a more efficient solution tailored to edge computing environments.
@@ II-A Traditional P2P Approaches
P2P architectures, such as Napster [10], BitTorrent [1], and IPFS [9], provide decentralized mechanisms where nodes act as both consumers and providers. While these systems excel in distributed replication and data sharing, they suffer from a critical drawback: the lack of bandwidth awareness, which results in suboptimal use of network resources, particularly in edge environments.
@@
To demonstrate this limitation, we conducted an experiment using Kraken [43]. As depicted on the left side of Fig. 1, our testbed consisted of two LANs connected via a 100 Mbps link. Each LAN included three Raspberry Pi (RPi) 4 Model B devices connected to a 1 Gbps switch. In LAN 1, two hosts seeded a large container image, while requests for the image originated from LAN 2. The experiment was repeated with 1, 2, and 3 local seeding peers in LAN 2. The results, shown in the right side of Fig. 1, highlight a significant inefficiency: despite having local peers available in LAN 2, approximately 10% of the data blocks were still fetched from remote peers in LAN 1, consuming over 95% of the transit network bandwidth. This behavior stems from BitTorrent’s peer selection strategy, which prioritizes global distribution over localized resource optimization. Consequently, the uplink bandwidth became saturated, negating the benefits of local peers and limiting the speed advantage of P2P systems in edge environments.
Tracker-based peer discovery is also vulnerable in dynamic edge environments given its centralized architecture. When a tracker-carrying node fails, it leads to fragmented networks, increased latency, and the requirement of manual intervention [12, 32, 25].
@@ II-B Image Structure-based Approaches
Another class of approaches, exemplified by Starlight [11], attempts to optimize container startup time by analyzing the structures of the requested images. Starlight removes unnecessary components and decouples the download and execution processes to minimize startup latency. While effective for some use cases, these optimizations rely on runtime trace collection, which assumes that future executions will follow similar patterns and the two phases can be effectively parallelized. This creates a critical drawback when dealing with diverse and unpredictable workloads. Moreover, Starlight retains a reliance on centralized registries, limiting its effectiveness in dynamic, resource-constrained edge environments.
Figure 2: Timeline diagram of a large model inference task. Top: Lazy Loading (e.g., Starlight [11]): (1) Download components (PyTorch, cuDNN library, etc.), (2) Download the model, (3) Load the ML framework, (4) Perform inference. Bottom: P2P (e.g., PeerSync): (1) Download the full image, (2) Load the ML framework, (3) Perform inference.
As shown in Fig. 2, Starlight’s lazy loading strategy prioritizes downloading smaller components (e.g., libraries) before larger files like machine learning models. However, this approach provides only minimal speedup for inference tasks, as the critical delay stems from downloading large model files before computation can begin. For example, uncompressed CUDA-accelerated versions of Meta’s Llama 3.1 [15] (e.g., ai/meta-llama:3.1-8B-Instruct-cuda-12.6) require approximately 21.4 GiB of disk space. This image is dominated by two major components: models (15 GiB or 70.10%, consisting of only four files using Safetensors [3]) and frameworks (6.2 GiB or 28.97%, comprising PyTorch [35] and CUDA libraries). Trace-based methods like Starlight struggle with such storage footprints, as the bulk of the data resides in a few large files. In contrast, P2P-based approaches, such as the one proposed in this work, optimize how data is transmitted rather than focusing on when or what to download. By leveraging multiple peers, P2P systems can accelerate the distribution of large data blocks, making them more effective for workloads dominated by large model components.
@@ II-C Insights
III System Design
III-A Overview
Fig. 3 illustrates the architecture of PeerSync, comprising four main components: Container API Interface (Sec. III-B), Download Engine (Sec. III-C), Embedded Tracker (Sec. III-D), and Cache Manager (Sec. III-E). Container API Interface provides a seamless and non-intrusive replacement for conventional Docker registries by implementing a suite of APIs. Download Engine coordinates the image retrieval process, combining traditional HTTP-based methods with innovative P2P-based approaches. Embedded Tracker maintains an autonomous and self-healing tracking system to ensure swarm integrity, eliminating reliance on centralized components. Cache Manager optimizes storage by managing cached images dynamically, prioritizing frequently used images and removing unused ones with minimal performance impact.
Figure 3: PeerSync architecture. The yellow boxes are PeerSync components.
Fig. 3 also depicts the lifecycle of an image within PeerSync, where red lines represent the retrieval process, and blue lines represent the removal process.
[…]
III-B Container API Interface
Container API Interface leverages the standardized APIs established by the Open Container Initiative (OCI). [^2] https://opencontainers.org/ Specifically, we adapt a subset of the Docker registry APIs to enable comprehensive container image retrieval and management while preserving simplicity for the user. This integration allows for seamless image handling, effectively masking the inherent complexity of the process. A critical aspect of this implementation is the management of image pulls by the Docker engine, which typically involves multiple HTTP requests: one to retrieve the image manifest and several more for each missing layer. The manifest, typically a JSON structure, contains essential details about the requested image, while layer requests return the actual data in the form of tarballs.
Our approach strategically optimizes these two API endpoints:
  • Manifest. Manifests are cached in an in-memory storage structure, which is continuously updated to reflect any upstream changes from Docker Hub.
  • Layer. This endpoint is a streamlined encapsulation that delegates the complexities of the multi-strategy image downloading process to Downloading Engine.
The API encapsulation layer within Docker Engine provides a simplified interface. Upon initialization, PeerSync queries the Docker Engine’s version via a UNIX domain socket and subsequently supplies runtime data to other system components. Essentially, Container API Interface acts as a bridge between Docker Engine and PeerSync’s internal mechanisms. Through this efficient encapsulation, PeerSync maintains precise control and monitoring of the Docker Engine, enabling a focus on optimizing internal operations.
III-C Downloading Engine
Downloading Engine serves as a core component in PeerSync, managing DHT network connections and facilitating popularity- and network-aware P2P downloading upon requests.
III-C1 Request Dispatcher
Request Dispatcher interfaces directly with Container API Interface (Sec. III-B). Upon receiving a download request, it evaluates the optimal strategy for image retrieval, deciding between downloads from a center registry and leveraging the P2P network. The process begins by assessing the size of the requested layer, using information stored in the manifest cache. Next, the Request Dispatcher determines whether the content is available within the nearby P2P network, taking into account time constraints during the search process.
P2P Downloading or Docker Hub? The efficiency of locating peers in the Distributed Hash Table (DHT) network is critical, as this process, which involves multiple network hops, can be time-consuming. Poor network conditions between peers, such as limited bandwidth or packet loss, can further extend the duration of peer discovery. In our context, such delays are unacceptable, as they negatively impact user experience by causing prolonged waiting times. To address this, we implement a time limit for peer aggregation. If image availability is confirmed within this timeframe, it indicates that image holders are present in nearby networks, making P2P downloading the preferred method. However, if the time limit expires without success, it suggests the absence of local image holders or their significant distance from the requesting node, prompting the system to switch to direct downloading from Docker Hub.

### v2

# PeerSync: Accelerating Containerized Model Inference at the Network Edge

@@ Abstract
[…]
@@
Edge computing, container image distribution, P2P architecture, local area network, model inference.
@@ I Introduction
Deploying a containerized application involves retrieving the necessary image layers from a container registry (e.g., Docker Hub), unpacking them, mounting them using a layered filesystem, and initiating the container entrypoint. These images, structured as a stack of layers representing filesystem segments, are stored and managed in centralized container registries. While these registries efficiently facilitate container image distribution in cloud environments, they encounter unique challenges in edge computing scenarios: (i) High latency: Distributed edge devices face substantial delays when fetching image layers over wide-area networks, leading to degraded responsiveness and increased cold-startup times for services [4, 9, 10]; (ii) Inefficiency under high concurrency: The non-parallel architecture of container registry uplinks often causes severe bottlenecks when handling concurrent image-pulling requests from multiple edge devices [10]. The growing size of AI models exacerbates this issue. To address these limitations, peer-to-peer (P2P) architectures have been explored to accelerate container image distribution while alleviating the load on centralized registries [4, 2, 11, 12, 13, 14, 8, 15, 7, 16]. P2P systems inherently offer distributed replication, parallel data transfers, and resilience against single points of failure. Although P2P-based solutions have shown promise in cloud environments, their effectiveness in edge settings is constrained by the unique characteristics of edge computing:
  • Limited network bandwidth. Unlike high-capacity data center networks, edge environments often has significantly constrained bandwidth, hindering fast data exchange [3].
  • Dynamic network topology. The dynamic nature of edge networks introduces frequent connectivity changes, requiring robust mechanisms to adaptively manage edge devices and maintain consistency without relying on centralized components such as static trackers [17].
@@
Table I: Comparison between PeerSync and state-of-the-art works. ‘F.D.’: ‘Fully Decentralized’, ‘S.D.’: ‘Semi Decentralized’.
Work PeerSync Kraken [16] Dragonfly [7] Starlight [10] EdgePier [9] FID [8]
@@
P2P flavor Impl. from scratch Impl. from scratch Impl. from scratch - IPFS BitTorrent
Scenario General Cloud Cloud WAN Edge General
Tracker placement dynamic static static - static static
While P2P-based solutions such as EdgePier [9], Kraken [16], and Starlight [10] address some of these issues and achieve great performance in most cases, they are still based on the traditional P2P architecture and consequently fall short in effectively balancing network efficiency, storage optimization, and adaptability to edge-specific conditions. More specifically, they lack fine-grained control mechanisms and rely on static trackers, making them unsuitable for fluctuating network conditions. Starlight [10] minimizes container startup latency by redesigning deployment protocols and storage formats but requires intrusive modifications to cloud registries, limiting its applicability. Moreover, existing solutions often neglect factors such as dynamic network quality, content popularity, and effective storage utilization, leaving room for improvement. For instance, our preliminary experiments (detailed in Sec. II) demonstrate that maintaining a single copy of any image layer within a local area network (LAN) can significantly enhance image-fetching speed without overloading local storage.
Crucially, the edge is not a minimized data center. Systems like Kraken [16] and Dragonfly [7], though highly optimized for homogeneous, stable cloud networks, assume persistent centralized trackers or super-nodes that become single points of failure under edge churn. EdgePier [9] improves decentralization but still presumes stable anchor nodes, which rarely exist in mobile or intermittently connected edge deployments. Meanwhile, proposals leveraging IPFS or BitTorrent [8] inherit protocol overheads ill-suited for low-resource devices. Even more recent learning-based approaches [19, 20, 21, 22], while theoretically appealing, require GPU/NPU resources, extensive training data, and stable feedback loops, rendering them impractical on typical edge hardware such as Raspberry Pi. To overcome these challenges, we propose PeerSync, a non-intrusive, P2P-based system tailored to the unique demands of the edge. PeerSync constructs a fully decentralized P2P network across different LANs. It features a P2P image download engine that optimizes distribution by periodically calculating the required content pieces for each missing layer of a requested image, leveraging a scoring system that evaluates both content popularity and the quality of the network connection between peers and the requesting edge device. Unlike traditional P2P architectures that rely on manually configured trackers, PeerSync autonomously elects trackers based on real-time metrics, including bandwidth availability and resource utilization, ensuring resilience and performance. PeerSync also employs a selective deletion mechanism to optimize storage utilization, maintaining only essential image layers within each LAN. We compare PeerSync with state-of-the-art solutions in Table I. In summary, our main contributions are as follows:
@@
    We design an edge container image distribution system PeerSync, featuring a high-performance downloading engine, fault-tolerant tracker with self-healing ability, and collaborative dynamic space reclamation for efficient caching.
@@
    PeerSync dynamically assigns scores to peers. Its scoring function jointly optimizes for (i) local network proximity (e.g., same subnet), (ii) real-time bandwidth stability, and (iii) layer popularity. This ensures that the majority of traffic remains within the LAN, directly addressing the high-latency and limited-bandwidth constraints of WAN links.
@@
    We implement PeerSync with 8000+ lines of Rust code, resulting in a statically linked binary of just 8.8 MB, making it easily deployable on edge devices. PeerSync is compatible with the Open Container Initiative (OCI) standard and integrates transparently with existing container runtimes.
@@
[…]
@@ II Motivation
In this section, we empirically and analytically dissect the limitations of two representative approaches: Kraken [16] and Starlight [10], motivating the need for a more efficient solution tailored to edge computing environments.
@@ II-A Traditional P2P Approaches
P2P architectures such as Napster [23], BitTorrent [24], and IPFS [25] enable decentralized content sharing by allowing nodes to act as both consumers and providers. While these systems excel in cloud or data center settings with abundant bandwidth and stable connectivity, they suffer from a critical flaw in edge environments: the lack of bandwidth awareness and network topology sensitivity. This leads to inefficient utilization of scarce inter-LAN links and undermines the very benefits P2P promises.
@@
To illustrate this, we conducted an experiment using Kraken [16], which builds upon BitTorrent’s protocol. As shown on the left side of Fig. 1, our testbed comprised two LANs connected via a 100 Mbps link, representative of typical edge-to-edge or edge-to-cloud uplinks. Each LAN included three Raspberry Pi 4 Model B devices (each with 4 GB RAM and 32 GB eMMC storage) connected through a 1 Gbps switch. In LAN 1, two hosts served as seeders for a large container image; download requests originated from LAN 2, with 1, 2, and 3 local peers progressively added as potential sources. The results (right side of Fig. 1) reveal a striking inefficiency: even when local peers were available in LAN 2, Kraken still fetched approximately 10% of the image blocks from remote peers in LAN 1. Crucially, this small fraction consumed over 95% of the inter-LAN bandwidth. This behavior stems from BitTorrent’s peer selection strategy, which prioritizes global peer diversity and piece availability over network locality. As a result, the narrow uplink becomes saturated, delaying downloads for all nodes and negating the latency advantage of local caching. Beyond bandwidth waste, P2P systems like Kraken impose significant storage pressure on edge devices. In a separate experiment, we distributed the top-10 most popular Docker Hub images across Kraken nodes. After downloading and decompressing these images, each node consumed 1408.54 MiB of disk space. For resource-constrained edge devices (e.g., those using SD cards or embedded flash with limited write endurance), such storage overhead is unsustainable, especially when images accumulate over time alongside sensor data or model checkpoints.
Furthermore, Kraken’s reliance on a centralized tracker for peer discovery introduces a single point of failure. In dynamic edge environments, where devices frequently join, leave, or experience intermittent connectivity, tracker outages lead to network fragmentation, prolonged peer discovery latency, and often require manual intervention to restore service [26, 27, 28]. This centralized coordination model is fundamentally at odds with the autonomy and resilience required in edge deployments.
@@ II-B Image Structure-based Approaches
An alternative strategy, exemplified by Starlight [10], seeks to minimize container startup latency by analyzing image structures and runtime behavior. Starlight removes unnecessary components (a process known as debloating [29, 30]) and decouples the download and execution phases through lazy loading: it prioritizes “hot” files (e.g., libraries) based on runtime traces collected during a profiling run. While effective for certain applications, this approach faces three critical limitations in edge AI scenarios:
  • Starlight assumes workload homogeneity and predictability. Starlight requires users to convert images into a custom format and execute them once to collect file-access traces. This profiling step assumes that future executions will follow similar patterns, which breaks down when deploying diverse or evolving AI workloads (e.g., switching from object detection to LLM inference). At scale, maintaining accurate traces for hundreds of edge applications becomes impractical.
  • Starlight offers minimal benefit for large-model inference tasks. As shown in Fig. 2, inference pipelines are typically dominated by two phases: (1) fetching the model and framework, and (2) computation. Crucially, the entire model must be available before inference can begin, as models are often stored as a few large files (e.g., in Safetensors [31] or PyTorch format [32]). Consider Meta’s Llama 3.1 [33] (ai/meta-llama:3.1-8B-Instruct-cuda-12.6). When uncompressed, this image occupies approximately 21.4 GiB of disk space, composed of only two major parts: Model weights: 15 GiB (70.10%), stored in just 4 files using Safetensors [31]; ML framework and dependencies: 6.2 GiB (28.97%), including PyTorch [32] and CUDA libraries. In such cases, downloading libraries first provides negligible speedup since the critical path remains the transfer of the 15 GiB model file. Trace-based methods like Starlight cannot accelerate this bottleneck because they optimize what and when to download, not how fast the data can be fetched.
Figure 2: Timeline diagram of a large model inference task. Top: Lazy Loading (e.g., Starlight [10]): Download components (PyTorch, cuDNN library, etc.) $\to$ Download the model $\to$ Load the ML framework $\to$ Perform inference. Bottom: P2P (e.g., PeerSync): Download the full image $\to$ Load the ML framework $\to$ Perform inference.
In addition, Starlight remains dependent on centralized registries. Despite its optimizations, it still pulls base images from Docker Hub or private registries, inheriting all the latency, bandwidth, and availability challenges of cloud-centric architectures. In disconnected or bandwidth-constrained edge settings, this dependency severely limits deployability.
@@ II-C Insights
III The PeerSync System
III-A Problem Formulation
We consider a distributed environment where nodes (including cloud instances and edge devices) pull container images from a central registry. Each image consists of a manifest and multiple immutable layers. When many nodes concurrently pull the same image, the registry becomes a bottleneck, leading to high latency and WAN bandwidth consumption. Let $\mathcal{P}$ be the set of peers, each holding a subset of layers $\mathcal{L}_{p}\subset\mathcal{L}$, where $p\in\mathcal{P}$. Peers are connected via heterogeneous, dynamic network links with unknown bandwidth and availability. A requesting peer $p_{\text{req}}$ must fetch all layers of a target image as quickly as possible, without prior knowledge of which peers have which layers or their current upload capacity. The problem is to minimize the end-to-end image pull time by opportunistically leveraging P2P transfers among peers, while operating without global coordination, runtime modifications, or assumptions about peer stability or connectivity.
III-B System Architecture Overview
PeerSync is designed as a drop-in enhancement for existing container ecosystems. As illustrated in Fig. 3, the system comprises four core components: API Interface, Downloading Engine, Embedded Tracker, and Cache Manager, which are orchestrated along two orthogonal threads: (i) the pull path, which handles on-demand layer retrieval, and (ii) the cache maintenance path, which runs continuously in the background to manage storage and peer metadata. The entire system exposes a standard OCI Distribution API, making it fully compatible with Docker, Podman, and other OCI-compliant clients. No modifications to the container runtime or user workflow are required; peers simply configure their client to use PeerSync as the registry endpoint.
Figure 3: The architecture of PeerSync.
A typical pull request begins when a container runtime issues an HTTP GET for a layer. The request is intercepted by the API Interface, which checks local cache availability. If the layer is absent, the request is forwarded to the Downloading Engine, where a lightweight Request Dispatcher decides whether to fetch the layer from upstream (via HTTP) or initiate a P2P transfer based on multiple factors. When P2P is selected, the engine queries the Embedded Tracker for candidate peers and coordinates block-level downloads. Upon successful retrieval, the layer is stored in the local Cache, and the response is returned to the client. Concurrently, the Cache Manager monitors global access patterns across layers and performs eviction based on a cost-aware policy that considers layer size, replication count within the LAN, and reuse likelihood. The Embedded Tracker, operating independently, maintains an up-to-date view of nearby peers through periodic gossip and self-election, ensuring coordination remains decentralized and resilient.
PeerSync attempts to join the P2P swarm during application startup. During bootstrap, PeerSync reads predefined nodes from the configuration and broadcasts bootstrap messages locally to join the swarm. If the bootstrap phase fails, PeerSync runs in a degraded mode as a proxy for the upstream registry and continues to bootstrap until a full P2P connection can be established. The predefined node list is dynamically updated with high-uptime peers, enabling PeerSync to withstand environmental changes.
III-C Component Design
III-C1 Container API Interface
Modern container ecosystems achieve remarkable interoperability through adherence to the OCI standards. The OCI defines both an Image Specification (governing image format and manifest structure) and a Distribution Specification (defining HTTP-based registry APIs for pulling images). To ensure PeerSync operates as a drop-in replacement without requiring any modification to existing runtimes or user workflows, it must fully comply with these standards. Crucially, the container runtime initiates image pulls via a well-defined sequence: one request for the image manifest (a JSON document listing layer digests and metadata), followed by one request per missing layer (each returning a gzipped tarball). Any deviation from this protocol would break compatibility. Thus, the primary motivation is to provide a transparent, standards-compliant facade that hides the internal P2P complexity while preserving the exact API contract expected by the runtime.
The Container API Interface implements a subset of the OCI Distribution Specification. It serves as the sole external entry point for the container runtime.
  • Manifest Handling. The manifest is typically small (often <5<5 KB) but frequently accessed and subject to upstream updates (e.g., latest tag). To ensure low-latency responses and immediate consistency, manifests are stored in an in-memory cache. This cache is kept up-to-date by periodically polling the upstream registry or by invalidating entries upon detecting a new pull request for a potentially stale tag.
  • Layer Handling. Layer requests are not served directly. Instead, the interface acts as a lightweight dispatcher, forwarding each blob request (identified by its digest) to the Downloading Engine (Sec. III-C). The engine is responsible for fulfilling the request, whether from local cache, a LAN peer, or the upstream registry, and streaming the data back through this interface.
Figure 4: The state machine of P2P downloading in PeerSync.
III-C2 Downloading Engine
A naive P2P approach, which attempts to find peers for every layer, fails to account for the empirical reality of container image composition. As shown in Table II, nearly half of all layers are smaller than 512 KiB, with a median size of just 1.03 MiB. For such small payloads, the latency of peer discovery in a Distributed Hash Table (DHT) or even a local multicast can exceed the time required to download the layer directly from a nearby registry. Conversely, large layers (e.g., base OS images) can benefit immensely from parallel, multi-source P2P downloads. The engine’s core motivation is to make a per-layer, context-aware decision that optimizes for both speed and system stability, dynamically choosing between P2P and direct upstream fetching.